# Mass and energy cascade in collisionless dark matter flow and relevant constraints on the dark matter particle massA.6

### v3

@@ 5.5 Particle random walk and halo density profile
$\begin{split}\frac{\partial P_{r}\left(\boldsymbol{X},t\right)}{\partial t}=D_{0}\frac{\partial}{\partial X_{i}}\left[\frac{\partial}{\partial X_{i}}\left(r^{2\gamma}P_{r}\left(\boldsymbol{X},t\right)\right)\right].\end{split}$ (76)
The corresponding solution of Eq. (76) in spherical coordinate is
$P_{r}\left(r,t\right)=\frac{(2-2\gamma)^{\frac{\gamma-2}{1-\gamma}}r^{-2\gamma}}{4\pi\left(D_{0}t\right)^{\frac{3-2\gamma}{2-2\gamma}}\Gamma\left(\frac{3-2\gamma}{2-2\gamma}\right)}\exp\left(-\frac{r^{2-2\gamma}}{4(1-\gamma)^{2}D_{0}t}\right).$ (77)
@@
Since the distribution function $P_{r}(r,t)$ is equivalent to the halo density, we find that halo density $\rho_{r}\propto r^{-2\gamma}$. From this insight, assuming that $\gamma$ is unknown, we can predict the value of $\gamma$ and recover the -4/3 law. This can be an independent check as an alternative approach other than deriving the -4/3 law from the scale-independent energy cascade $\varepsilon$ (see Eqs. (69) and (70)):
@@
Since the waiting time $\tau_{gr}\propto\Phi(r)^{-1}\propto r^{-\gamma}$ (Eq. (55)), the halo density should scale as $\rho_{r}\propto r^{-2\gamma}$ from Eq. (77). The halo mass enclosed in $r$ scales as $m_{r}\propto\rho_{r}r^{3}\propto r^{3-2\gamma}$. The local potential at $r$ should scale as $\Phi(r)\propto Gm_{r}/r\propto r^{3-2\gamma-1}$. The waiting time of the particle at $r$ satisfies $\tau_{gr}\propto\Phi^{-1}\propto r^{2\gamma-2}$ (from Eq. (55)). Combined with $\tau_{gr}\propto r^{-\gamma}$, we should have $2\gamma-2=-\gamma$ such that $\gamma=2/3$ and the inner density slope should be $32\gamma=4/$ 3. Therefore, the halo density scales as $\rho_{r}\propto r^{-4/3}$ and the waiting time $\tau_{gr}\propto r^{-2/3}$. It should be noted that the random walk theory for halo formation and evolution confirms the -4/3 law ($\rho_{r}\propto r^{-4/3}$) that is also predicted by the energy cascade in haloes with a scale-independent rate (Eq. (70)). The predictions of the -4/3 scaling law are tested against the simulations in Figs. 13 and 14.
@@
Similarly to the halo mass function (Eq. (33)), the exponent $\gamma$ can be different in two different ranges of scales, the inner halo below the scale radius $r_{s}$ and the outer halo above $r_{s}$. Using two different $\gamma$ for the $r$ dependence of waiting time $\tau_{g}(r)\propto r^{-\gamma}$, that is, $\gamma_{1}$ and $\gamma_{2}$ for two different ranges, the double-$\gamma$ distribution function can be obtained based on the single-$\gamma$ distribution in Eq. (77)
$P_{r}\left(r,t\right)=\frac{(2-2\gamma_{2})^{\frac{2\gamma_{1}-2-\gamma_{2}}{1-\gamma_{2}}}r^{-2\gamma_{1}}}{4\pi\left(D_{0}t\right)^{\frac{3-2\gamma_{1}}{2-2\gamma_{2}}}\Gamma\left(\frac{3-2\gamma_{1}}{2-2\gamma_{2}}\right)}\exp\left(-\frac{r^{2-2\gamma_{2}}}{4(1-\gamma_{2})^{2}D_{0}t}\right).$ (78)
@@
Introducing the conventional scale radius $r_{s}(t)$ where the logarithmic slope of $P_{r}(r,t)$ equals -2, we should have
$4(1-\gamma_{2})^{2}D_{0}t=\frac{2-2\gamma_{2}}{2-2\gamma_{1}}r_{s}^{2-2\gamma_{2}}.$ (79)
@@
The time variation of the distribution function is absorbed into the scale radius $r_{s}(t)$. The double-$\gamma$ distribution (Eq. (81)) reduces to the Einasto profile with $\alpha=2\beta$ or $\gamma_{1}=0$. The cumulative distribution in spherical coordinates reads

### v4

@@ 4.1 Particle random walk and distribution in haloes
where $D_{0}(t)$ is a proportional constant. The closer to the halo center (smaller $r$), the smaller the diffusivity or the longer the waiting time, and the higher the density of the particles. By introducing a time $\tau$ proportional to the scale factor $\tau=(3/2)at_{0}$ ($t_{0}$ is the present epoch), $a^{-1/2}$ can be absorbed, and the original Eq. (53) reads
$\frac{d\boldsymbol{X}_{\tau}}{d\tau}=\sqrt{2D_{P}(\boldsymbol{X}_{\tau})}\boldsymbol{\xi}\left(\tau\right).$ (55)
@@
$\begin{split}\frac{\partial P_{r}\left(\boldsymbol{X},\tau\right)}{\partial\tau}=D_{0}\frac{\partial}{\partial X_{i}}\left[\frac{\partial}{\partial X_{i}}\left(r^{2\gamma}P_{r}\left(\boldsymbol{X},\tau\right)\right)\right].\end{split}$ (56)
The corresponding solution of (56) in spherical coordinates reads
$P_{r}\left(r,\tau\right)=\frac{(2-2\gamma)^{\frac{\gamma-2}{1-\gamma}}r^{-2\gamma}}{4\pi\left(D_{0}\tau\right)^{\frac{3-2\gamma}{2-2\gamma}}\Gamma\left(\frac{3-2\gamma}{2-2\gamma}\right)}\exp\left(-\frac{r^{2-2\gamma}}{4(1-\gamma)^{2}D_{0}\tau}\right).$ (57)
@@
Since the distribution function $P_{r}(r,\tau)$ is equivalent to the halo density, we find that the halo density $\rho_{r}\propto r^{-2\gamma}$ for the waiting time $\tau_{gr}(r)\propto r^{-\gamma}$. From this insight, assuming that $\gamma$ is unknown, we can predict $\gamma=2/3$ for gravity-dominant halo evolution.
@@
Since the waiting time $\tau_{gr}\propto\Phi(r)^{-1}\propto r^{-\gamma}$, the halo density should scale as $\rho_{r}\propto r^{-2\gamma}$ from Eq. (57). The halo mass enclosed in $r$ scales as $m_{r}\propto\rho_{r}r^{3}\propto r^{3-2\gamma}$. The local potential at $r$ should scale as $\Phi(r)\propto Gm_{r}/r\propto r^{3-2\gamma-1}$. The waiting time of the particle at $r$ satisfies $\tau_{gr}\propto\Phi^{-1}\propto r^{2\gamma-2}$. Combined with $\tau_{gr}\propto r^{-\gamma}$, we have
$2\gamma-2=-\gamma\quad\textrm{such that}\quad\gamma=2/3.$ (58)
The density slope should be $2\gamma=4/3$. Therefore, the halo density scales as $\rho_{r}\propto r^{-4/3}$ and the waiting time $\tau_{gr}\propto r^{-2/3}$. This interesting result demonstrates that in the absence of the effect of the initial spectrum, gravity should drive the haloes toward a limiting slope of 4/3-4/3. For a realistic CDM spectrum, although the core structure can be affected by the initial spectrum, the universal slope of 4/3-4/3 should be true for large halos of characteristic mass $m_{h}^{*}$ or for haloes on scales close to $r_{s}$. It should be noted that the random walk theory for halo formation and evolution confirms a -4/3 law ($\rho_{r}\propto r^{-4/3}$) that can also be predicted by the energy cascade in haloes with a scale-independent rate (Eq. (71)). The predictions of the -4/3 scaling law are tested against the simulations in Figs. 11 and 12.
@@
Similarly to the halo mass function (Eq. (26)), the exponent $\gamma$ can be different in two different scale ranges, the inner halo below the scale radius $r_{s}$ that is affected by the initial spectrum and the outer region around scale $r_{s}$ dominated by gravitational dynamics. Using two different $\gamma$, that is, $\gamma_{1}$ and $\gamma_{2}$ for two different ranges, the double-$\gamma$ distribution function can be obtained based on the single-$\gamma$ distribution in Eq. (57),
$P_{r}\left(r\right)=\frac{(2-2\gamma_{2})^{\frac{2\gamma_{1}-2-\gamma_{2}}{1-\gamma_{2}}}r^{-2\gamma_{1}}}{4\pi\left(D_{0}\tau\right)^{\frac{3-2\gamma_{1}}{2-2\gamma_{2}}}\Gamma\left(\frac{3-2\gamma_{1}}{2-2\gamma_{2}}\right)}\exp\left(-\frac{r^{2-2\gamma_{2}}}{4(1-\gamma_{2})^{2}D_{0}\tau}\right).$ (59)
@@
Introducing the conventional scale radius $r_{s}(t)$ where the logarithmic slope of $P_{r}(r,t)$ is equal to -2, we should have
$4(1-\gamma_{2})^{2}D_{0}\tau=\frac{2-2\gamma_{2}}{2-2\gamma_{1}}r_{s}^{2-2\gamma_{2}}.$ (60)
[…]
@@
Both parameters $\alpha$ and $\beta$ can be related to $\gamma$ for the scale dependence of waiting time $\tau_{gr}(r)\propto r^{-\gamma}$. The time variation of the distribution function is absorbed into the scale radius $r_{s}(t)$. The double-$\gamma$ distribution (Eq. (62)) reduces to the Einasto profile with $\alpha=2\beta$ or $\gamma_{1}=0$. The cumulative distribution in spherical coordinates reads
$\int_{0}^{x}P_{r}\left(y\right)4\pi{y}^{2}d{y}=1-\frac{\Gamma\left(\frac{1}{\alpha}+\frac{1}{\beta},\frac{x^{\alpha}}{\beta}\right)}{\Gamma\left(\frac{1}{\alpha}+\frac{1}{\beta}\right)},$ (64)
where $\Gamma(x,y)$ is an upper incomplete gamma function.
4.2 Halo density profiles and universal scaling laws
The halo density profile can be inferred from the particle radial distribution function $P_{r}$. The well-known core–cusp problem highlights the tension between the cuspy inner profiles produced by cosmological cold dark matter (CDM)–only simulations and the approximately constant-density cores indicated by observations. In CDM simulations, the inner density typically follows $\rho\propto r^{s}$, with reported logarithmic slopes s in the range -1.0 to -1.5 (see Introduction Section 1). There is neither consensus on a single value of s nor a robust first-principles theory that uniquely predicts it. This non-universality reflects the joint influence of the primordial power spectrum and nonlinear gravitational dynamics on halo structure: the inner halo retains greater memory of the initial spectrum and thus exhibits CDM-dependent variation in $s$, whereas the region near the scale radius $r_{s}$ is more strongly shaped by gravity and correspondingly displays a more universal profile.
A double-$\gamma$ density profile can accommodate both effects through two distinct parameters. We interpret $\gamma_{1}$ as an effective descriptor of the inner structure, governed by the primordial power spectrum, whereas $\gamma_{2}$ encapsulates the influence of gravitational dynamics. The general expression reads
$\rho_{r}(r,t)=\rho_{s}(m_{h},t)\left(\frac{r}{r_{s}}\right)^{\frac{\alpha}{\beta}-2}\exp\left(\frac{1}{\beta}\left(1-{\left(\frac{r}{r_{s}}\right)^{\alpha}}\right)\right),$ (65)
where $\rho_{s}(t)$ is the density at scale radius $r_{s}(t)$. This four-parameter double-$\gamma$ density profile ($\rho_{s}$, $r_{s}$, $\alpha$, and $\beta$ in Eq. (65)) reduces to the three-parameter Einasto profile $($$\rho_{s}$, $r_{s}$, and $\alpha$) with $\alpha=2\beta$.
For massive haloes near the characteristic mass $m_{h}^{*}$, the structure across most spatial scales is governed primarily by gravitational dynamics and the associated radial mass and energy flux(Xu:2023-Universal-scaling-laws-and-density-slope; Xu:2023-Dark-matter-halo-mass-functions-and), producing a universal logarithmic density slope of -4/3. The same behavior also holds in the vicinity of the scale $r_{s}$ for haloes of arbitrary mass $m_{h}$. Under these conditions, the double-$\gamma$ density profile is obtained with $\alpha/\beta=2/3$ in Eq. (65), or $\gamma_{1}=2/3$, and it reads:
where $\beta_{r}\approx 1$ is an amplitude parameter, $\beta$ is a shape parameter in Eq. (63). Here, we introduce a parameter $\varepsilon$ (m2s-3) to replace the density $\rho_{s}(t)$ in Eq. (65) to respect the -4/3 law on scales near $r_{s}$. This -4/3 law for density should only involve the gravitational constant $G$ that reflects the dynamics governed by gravity (see the derivation of scaling laws in Eq. (71)):
$\begin{split}&\rho_{s}(m_{h},t)=\beta_{r}\exp(-1/\beta)\varepsilon^{2/3}G^{-1}r_{s}^{-4/3}.\end{split}$ (67)
The parameter $\varepsilon(m_{h},z)$ represents the rate of the energy flow in haloes of mass $m_{h}$ (defined in Eq. (70)), while $\varepsilon_{u}=\varepsilon(m_{h}^{*},z)\approx 10^{-7}m^{2}/s^{3}$ is the rate of the energy flow in haloes with characteristic mass $m_{h}^{*}(z)$. This is also the rate of "dissipation" for global-scale cosmic energy (Section B). Both parameters and the associated mass and energy flow in nonequilibrium haloes control the halo structure near $r_{s}$. Its physical meaning is discussed in the next subsection 4.3. The values are estimated from N-body simulations (Figs. 14 and 15).
In particular, for small $r$, the inner density for gravitational dynamics dominant haloes (Eq . (66)) is reduced to the -4/3 law:
$\rho_{r}(r,m_{h},z)=\beta_{r}\varepsilon^{\frac{2}{3}}G^{-1}r^{-\frac{4}{3}}\propto m_{h}^{\frac{4}{9}}a^{-\frac{2}{3}}r^{-\frac{4}{3}}\quad\textrm{for}\quad r\rightarrow 0.$ (68)
The redshift dependence is absorbed into the parameter $\varepsilon$. For haloes of characteristic mass $m_{h}^{*}(z)\propto a^{3/2}$, there exists a small-scale permanence for the halo density, i.e., the density profiles at different redshifts $z$ converge to a time-unvarying scaling,
The scaling law shows the density of the halo $\rho_{r}(r)\propto m_{h}^{4/9}a^{-2/3}$ for haloes of mass $m_{h}$. Clearly, halo mass accretion leads to an increase in $m_{h}$ and density $\rho_{r}$, while expansion leads to a decrease in $\rho_{r}$ over time. The two competing effects are equal for haloes with characteristic mass $m_{h}^{*}\propto a^{3/2}$ such that the inner density $\rho_{r}$ for haloes with characteristic mass $m_{h}^{*}$ remains constant over time. This is the so-called small-scale permanence for the halo density. The density of haloes of mass $m_{h}^{*}$ at different redshifts collapses into the same -4/3 law. For old and small haloes with a mass less than $m_{h}^{*}$, the mass accretion is slower with a smaller $\varepsilon$, leading to a decrease in density $\rho_{r}$ over cosmic time $t$, as shown in Eq. (68).
To validate the predicted halo density and scaling laws, we present the N-body simulation results. Figure 11 demonstrates the small-scale permanence for the halo density profile, similarly to the small-scale permanence for the halo group mass $m_{g}$ (Fig. 9). The figure plots the time evolution of the average density profiles for all haloes with characteristic masses $m_{h}^{*}(z)$ from Illustris simulations at different redshifts. All haloes have a limiting constant density slope of -4/3. The density profiles of all dark matter haloes of mass $m_{h}^{*}$ converge to the predicted time-unvarying scaling (solid red line from Eq. (69)) on small scales, i.e., the small-scale permanence. The -4/3 scaling law in Fig. 11 should extend to smaller and smaller scales until the smallest scales. That scale can be the free streaming scale, which is dependent on the nature and mass of dark matter particles. The properties of dark matter particles may be obtained by extending these scaling laws to the smallest scale (Xu:2022-Postulating-dark-matter-partic).
Figure 12 presents the density profiles from Illustris simulations for (composite) haloes of different masses $m_{h}$ at $z=0$. The solid lines show the average halo density profiles for all haloes with a mass between $10^{\pm 0.1}m_{h}$ from the Illustris simulation, where the halo mass $m_{h}$ is between $10^{8}M_{\odot}$ and $10^{14}M_{\odot}$. The thick dashed (straight) line represents the -4/3 scaling laws for haloes of mass $m_{h}$ that involve the rate of cascade $\varepsilon\propto m_{h}^{2/3}$ (Eq. (94)). Here, $\varepsilon=\varepsilon_{u}$ for haloes of characteristic mass $m_{h}^{*}$ (Eq. (94)). The figure demonstrates the -4/3 law that we obtained from the energy cascade in haloes. That scaling was in agreement with the density profiles of haloes of more than six orders of magnitude. The thin dashed lines in both Figs. 12 and 11 present the double-$\gamma$ density profiles from Eq. (66) that agree well with the simulation data.
Universal scaling laws are well known to be associated with the energy cascade in fluid turbulence (Richardson:1922-Weather-Prediction-by-Numerica; Kolmogoroff:1941-Dissipation-of-energy-in-the-l). Next, we focus on the scaling laws in spherical haloes governed by gravitational dynamics, which can be derived from the particle distribution. Because of the slope $\rho_{r}\propto r^{-4/3}$ (Eq .(57)), we have the total mass $m_{r}\propto\rho_{r}r^{3}\propto r^{5/3}$ enclosed within r, the characteristic velocity $v_{r}^{2}\propto Gm_{r}/r\propto r^{2/3}$, and the time $t_{r}\propto r/v_{r}\propto r^{2/3}$ on scale r. Next, a simple definition of $\varepsilon$ that is independent of scale r reads,
The microscopic interpretation of $\varepsilon$ is associated with the energy transfer in haloes in the next section (Eq. (79)).
The rate of the energy cascade $\varepsilon$ is independent of $r$ on a certain range of scales. The 2/3 law can be easily obtained for kinetic energy $v_{r}^{2}\propto(-\varepsilon r)^{2/3}$ from Eq. (70). With the virial theorem ($v_{r}^{2}\propto Gm_{r}/r$), the 5/3 scaling law can be recovered for the mass enclosed within $r$, i.e., $m_{r}\propto r^{5/3}$. We can write the scaling laws for mass $m_{r}$, density $\rho_{r}$, velocity $v_{r}$, time $t_{r}$, and kinetic energy $v_{r}^{2}$, and mass flux $\dot{m}_{r}$ on scale $r$, all determined by three and only three quantities $\varepsilon$, $G$, and scale $r$ reflecting gravity-dominant dynamics:
$\begin{split}&m_{r}=\alpha_{r}\varepsilon^{2/3}G^{-1}r^{5/3}\textrm{,}\quad\rho_{r}=\beta_{r}\varepsilon^{2/3}G^{-1}r^{-4/3},\\ &v_{r}^{2}=\beta_{h}^{-2/3}(\varepsilon r)^{2/3}\textrm{,}\quad t_{r}=0.5\beta_{h}^{-2/3}\varepsilon^{-1/3}r^{2/3},\\ &v_{r}^{2}=\beta_{h}^{-2/3}\alpha_{r}^{-2/5}(\varepsilon Gm_{r})^{2/5},\quad\dot{m}_{r}=4\pi\beta_{h}^{-1/3}\beta_{r}\varepsilon G^{-1}r\end{split}$ (71)
where $\alpha_{r}$ and $\beta_{r}$ are two numerical factors that can be determined by fitting the data (Fig. 11). The evolution of these quantities is absorbed into $\varepsilon(m_{h},z)$. The relations between numerical factors are read as
$\begin{split}\beta_{h}=\frac{1}{3},\quad\gamma_{r}\alpha_{r}=\beta_{h}^{-2/3},\quad\textrm{and}\quad\beta_{r}=\frac{5}{4\pi}\alpha_{r}\beta_{h}\end{split}$ (72)
Another approach to derive these scaling laws using the dimensional argument is relatively simpler but more heuristic. When a statistically steady state is established due to gravitational dynamics, the fast motion on small scales does not feel the slow motion on large scales directly, except through the scale-independent rate of the energy cascade $\varepsilon$. The flow fields on these scales are statistically similar so that all relevant physical quantities can be determined by and only by three quantities: $\varepsilon$, $G$, and scale $r$. Therefore, the scaling laws for any quantity $Q=\varepsilon_{u}^{x}G^{y}r^{z}$ can be determined from the dimensional analysis to obtain the exponents $x$, $y$, and $z$. This approach leads to the same scaling laws as in Eq. (71).
4.3 Particle random walk and energy cascade
Fully virialized haloes cannot sustain mass or energy transport in the limit of vanishing energy flux ($\varepsilon$=0); Therefore, $\varepsilon$=0 represents the formal equilibrium state toward which haloes asymptotically relax but never truly attain. In this section, we examine the mass and energy fluxes in nonequilibrium haloes that are evolving toward that limiting equilibrium. These fluxes are driven by the stochastic random walk of individual particles within the halo potential, which leads to a transfer of mass and energy across scales. This microscopic diffusion of particles is analogous, on the internal halo scale, to the random walk of haloes in the mass space that underlies the mass and energy cascade in the mass space (Section 3).